when the son was 8 years old the father was 31 years old. Now the father is 2 times older than the son. How old is the son, and how old is the father
Found 2 solutions by richwmiller, TamiSrcKa:
Answer by richwmiller(17219)

huh? I hope you mean the father is twice the age of the son.
f=31,
s=8,
2*(8+x)=31+x
16+2x=31+x
x=15
15 years later, the son is 23 and the father is 46


Answer by TamiSrcKa(4)

before
father 31
son 8
.....
now
father 2x
son x
____________
Solution:
31-8=2x-x
23=x
>>> The son is 23, and the father is 2x=2*23=46.